### v1

@@ 4.1 CyberBattleSim Environments
CyberBattle Capture-the-Flag (CTF) Environment. This environment uses a loosely connected, hub-and-spoke topology with nodes representing services like web apps, GitHub, Azure storage, SharePoint, and cloud VMs. Each node supports multiple exploit paths, such as scanning content, mining Git history, or reusing leaked credentials. Supported actions include SearchEdgeHistory, ScanPageContent, NavigateWebDirectory, AccessDataWithSASToken, and privilege escalations. Vulnerabilities are contextually embedded (e.g., exposed tokens, weak MySQL/SSH logins), requiring chained reasoning. Its non-linear structure supports analysis of both strategic variability and tactical adaptation.
@@
CyberBattleChain Environment. CyberBattleChain simulates a linear, fixed-sequence network topology alternating between Linux and Windows nodes: start → (Linux → Windows)n → Linux[Flag]. Progression requires exploiting platform-specific vulnerabilities and reusing credentials revealed upon node compromise (e.g., SSH credentials from Windows to the subsequent Linux node). Embedded traps lure agents into suboptimal paths, requiring strategic discernment to avoid costly dead ends. The environment ends with a high-value target ("flag" node), making it ideal for evaluating long-horizon planning, sequential decision-making, and explainability in high-stakes attacker behaviour.
5 Multi-Layer Explainability Framework for RL-Based Attackers
@@
To address the opacity of RL agents in adversarial environments, we propose a unified, multi-layer explainability framework that reveals both strategic intent and tactical decision-making over time. Grounded in core explainable AI principles, such as transparency, temporal insight, and actionable reasoning, our framework operates at two complementary levels:
@@
    Strategic-Level (MDP): Modelling attacker behaviour under uncertainty, capturing exploration-exploitation dynamics and phase transitions across the attack lifecycle.
@@
    Tactical-Level (Policy): Tracing evolving action preferences and identifying high-impact decision points using temporal Q-value analysis and prioritised experience replay.
@@
5.1 Strategic-Level (MDP) Explainability
We formalise attacker behaviour as Partially Observable Markov Decision Process, defined by the tuple $(S,A,\Omega,O,T,R,\gamma)$, where $S$ is the set of true (hidden) states, $A$ the action space, $\Omega$ the set of possible observations, $O:S\times A\rightarrow\mathcal{P}(\Omega)$ is the observation function (a probability distribution over observations), $T:S\times A\rightarrow\mathcal{P}(S)$ the state transition function, $R:S\times A\rightarrow\mathbb{R}$ the reward function, and $\gamma\in[0,1]$ the discount factor. This formulation captures uncertainty and limited visibility adversaries face in real networks, and allows us to analyse strategic adaptations over time.
5.1.1 Exploration-Exploitation Dynamics
@@
5.1.2 Phase Transitions and Behaviour Evolution
@@
    Early Phase: In the early stages of an attack, strong defences and limited visibility are expected to constrain the agent’s progress, prompting a cautious, reconnaissance-driven approach. This phase is typically characterised by uncertainty, minimal network compromise, and broad probing, as the agent gathers information and identifies potential vulnerabilities. To capture this behaviour, our framework isolates transitions from early phases of the attack, where only a small fraction of the network has been compromised, enabling analysis of how strategies form under high uncertainty ($C_{t}<\text{Threshold}$).
@@
5.2 Tactical-Level (Policy) Explainability
@@
5.2.1 Temporal Q-Value Evolution and Action Preference
We track the evolution of Q-values over time to analyse how the agent’s assessment of state-action pairs changes with experience. In RL settings, Q-values begin largely undifferentiated due to sparse early feedback, and progressively evolve to reflect long-term reward expectations. Our framework leverages this temporal progression to surface how tactical preferences may emerge over time, offering insight into the agent’s evolving prioritisation of specific actions. The Q-values are updated via the Q-learning rule [21]:
@@ 5.2.1 Temporal Q-Value Evolution and Action Preference
5.2.2 Policy Interpretation via Prioritised Experience Replay
@@
By integrating strategic-level progression with tactical decision analysis, our framework provides a comprehensive and temporally grounded explanation of RL-based attacker behaviour. It elucidates not only the actions taken by the agent, but also the underlying rationale, timing, and policy evolution driving those decisions. This layered interpretability enhances transparency, supports informed defensive planning, and enables the development of phase-aware countermeasures aligned with the agent’s behavioural dynamics.
Table 1: Description of CyberBattleSim Environment.
@@
6 Experimental Evaluation
We evaluate our framework across five experimental setups designed to highlight key explainability dimensions over raw performance: (1) baseline policy selection, (2) exploration-exploitation behaviour, (3) phase-aware behavioural shifts, (4) temporal Q-value evolution, and (5) strategic learning signals via PER. These experiments serve as illustrative case studies demonstrating how our framework can analyse and interpret RL agents in adversarial cybersecurity scenarios.
Environment Description. We evaluated our framework using CyberBattleSim[18] across two core environments: (1) CTF, a 12-node targeted attack scenario, and (2) CyberBattleChain, with three scalable variants, CC22, CC100, and CC500, simulating networks of 22, 100, and 500 nodes, respectively. Table 1 summarises these configurations, where “Size” denotes exploitable nodes and “Max Nodes” includes the entire infrastructure, such as firewalls and auxiliary components.
@@ 6 Experimental Evaluation
Figure 1: Cumulative Rewards Comparison of Attacker Policies Across Environments (The shaded region in the figures represents the standard deviation of cumulative rewards across training steps for each agent).
6.1 Setup 1: Baseline Policy Selection
Results. Figure 1 presents a comparative summary of agent performance across all environments and the results show that DQL significantly outperformed the baseline agents in terms of stability, scalability, and cumulative reward. Classical methods such as Tabular Q-Learning and Credential Lookup struggled in larger environments (CC100+, with no viable results in CC500). Random search underperformed in all structured settings. While the Exploiting DQL variant achieved slightly higher cumulative rewards, it is not a standalone method; it relies on a fully trained DQL agent and simply exploits the final policy without any further training or adaptation. As such, it does not reflect the ongoing decision-making process that our explainability framework aims to explain. We therefore, selected DQL as the attacker policy for all subsequent experiments. We emphasise that the results in Figure 1 serve to establish a realistic setting to demonstrate the utility of our multi-layer explainability framework.
6.2 Setup 2: Exploration-Exploitation Dynamics
To analyse how exploration-exploitation strategies shape agent behaviour, we compare two strategies: Early and Standard Exploration and track their impact on cumulative reward and node discovery rate across environments. This reveals whether agent behave opportunistically, methodically, or inefficiently over time.
Results. For Early Exploitation, we initialised $\epsilon$ at 0.001 and gradually increased it to 0.9 at a growth rate of 0.005. In contrast, Standard Exploration began with $\epsilon=0.9$ and decayed to 0.01 at a rate of 0.95. Training was conducted for 30 episodes (200 iterations) in CTF and CC22, 40 episodes (300 iterations) in CC100 and 100 episodes (300 iterations) in CC500. Figure 2 shows that in smaller topologies (CTF and CC22), Standard Exploration quickly outperformed Early Exploitation, benefitting from more aggressive probing. In larger environments (CC100 and CC500), this advantage widened considerably, with Early Exploitation often stagnating at lower cumulative rewards, especially in CC500, where Standard Exploration achieved a notable reward surge after sustained, wide-ranging probing. These patterns are echoed in the node discovery rates in Figure 3 where Standard Exploration uncovered more nodes earlier, while Early Exploitation remained constrained.
Our findings highlight how exploration strategy shapes agent performance and behavioural style. Standard Exploration encourages deliberate, reconnaissance-driven tactics, while Early Exploitation creates opportunistic agents well-suited to complex environments. Our framework surfaces these differences, supporting interpretable assessments of attacker risk posture and planning style.
@@ 6.2 Setup 2: Exploration-Exploitation Dynamics
Figure 2: Impact of Exploration Strategies on Cumulative Rewards Across Environments.
@@
Figure 3: Impact of Exploration Strategies on Node Discovery Rate Across Environments.
@@
Figure 4: Cumulative Rewards Comparison in Early vs. Late Phase Attack Across Environments.
6.3 Setup 3: Phase-Aware Behaviour Evolution
This setup supports phase-aware explainability by segmenting attacker behaviour into early and late stages, using the compromise ratio $C_{t}$ (the proportion of compromised nodes) to highlight when strategy shifts occur.
Results. Agents were trained for 20 episodes (200 iterations) in CTF and CC22 environments, 100 episodes (400 iterations) in CC100, and 200 episodes (400 iterations) in CC500. Early phase is defined as network compromise ratio $C_{t}<0.5$, and late phase as $C_{t}\geq 0.5$. Figure 4 shows that early-phase rewards remain modest due to limited visibility and cautious probing. Late-phase curves begin with higher rewards and steeper slopes, reflecting the agent’s growing confidence and accumulated access.
In CyberBattleChain environments, the attack path is fixed: each node grants access to the next. Behavioural variation therefore reflects how quickly and efficiently the agent escalates privileges, selects exploits, and avoids traps, rather than differences in path selection. The observed late-phase deceleration likely reflects structural limitations rather than a shift in strategy.
In contrast, the CTF environment offers a more open topology. Here, phase-aware analysis can uncover how strategy evolves across multiple paths, revealing, for example, which nodes are prioritised once situational awareness improves. These insights can guide defenders in identifying high-risk services and timing of potential escalation.
More generally, in flexible or real-world networks, this type of analysis could help detect when an agent shifts from reconnaissance to exploitation, supporting adaptive defences or deception deployment. For developers, reward trajectories across phases provide feedback on convergence quality and policy generalisation. This setup thus enables interpretable comparisons across environments with varying structural complexity.
@@ 6.3 Setup 3: Phase-Aware Behaviour Evolution
Figure 5: Emergence of Action Preferences via State-Aggregated Q-Values Across Episodes.
6.4 Setup 4: Temporal Q-Value Analysis

### v2

@@ IV-A CyberBattleSim Environments
CyberBattle Capture-the-Flag (CTF) Environment. This environment uses a loosely connected hub-and-spoke topology, with nodes representing enterprise services like web applications, GitHub, Azure storage, SharePoint, and cloud VMs. Each node exposes multiple context-dependent exploit paths (e.g., scanning content, mining Git history, credential reuse), requiring chained reasoning to progress. Supported actions include SearchEdgeHistory, ScanPageContent, NavigateWebDirectory, AccessDataWithSASToken, and local and remote privilege escalation actions, with vulnerabilities embedded through exposed tokens, weak authentication, and misconfigurations. The environment’s non-linear structure supports analysis of both strategic variation and tactical adaptation.
Figure 1 illustrates the agent’s discovery and control escalation in the ToyCTF environment. Early episodes show limited visibility and control of only the client node (Figure 1(a)). As training progresses (Figures 1(b)–(d)), the agent uncovers adjacent services and additional assets (e.g., Website, GitHubProject, AzureStorage), leveraging lateral movement and remote exploits. In later stages (Figures 1(e)–(f)), the agent attains broad visibility and compromises critical assets such as AzureResourceManager and AzureVM, demonstrating the transition from initial access to network-wide dominance and highlighting ToyCTF’s suitability for analysing enterprise-scale attack progression.
@@
CyberBattleChain Environment. CyberBattleChain simulates a linear, fixed-sequence network topology alternating between Linux and Windows nodes: start → (Linux → Windows)n → Linux[Flag]. Progression requires exploiting platform-specific vulnerabilities and reusing credentials revealed upon node compromise (e.g., SSH credentials from Windows to the subsequent Linux node). Embedded traps lure agents into suboptimal paths, requiring strategic discernment to avoid costly dead ends. The environment ends with a high-value target (“flag” node), making it ideal for evaluating long-horizon planning, sequential decision-making, and explainability in high-stakes attacker behaviour.
V Multi-Layer Explainability Framework for RL-Based Attackers
@@
To address the opacity of RL agents in adversarial environments, we propose a unified, multi-layer explainability framework that reveals both strategic intent and tactical decision-making over time. Grounded in core explainable AI principles, such as transparency, temporal insight, and behavioural attribution, our framework operates at two complementary levels:
@@
    Strategic-Level (MDP): Models attacker progression through exploration–exploitation dynamics and phase transitions over the attack lifecycle.
@@
    Tactical-Level (Policy): Traces evolving action preferences and learning signals using temporal Q-value analysis and Prioritised Experience Replay (PER).
@@
V-A Strategic-Level (MDP) Explainability
We formalise attacker behaviour as a Partially Observable Markov Decision Process, defined by the tuple $(S,A,\Omega,O,T,R,\gamma)$, where $S$ is the set of true (hidden) states, $A$ the action space, $\Omega$ the set of possible observations, $O:S\times A\rightarrow\mathcal{P}(\Omega)$ is the observation function (a probability distribution over observations), $T:S\times A\rightarrow\mathcal{P}(S)$ the state transition function, $R:S\times A\rightarrow\mathbb{R}$ the reward function, and $\gamma\in[0,1]$ the discount factor. This formulation captures uncertainty and limited visibility adversaries face in real networks, and allows us to analyse strategic adaptations over time.
V-A1 Exploration-Exploitation Dynamics
@@
V-A2 Phase Transitions and Behaviour Evolution
@@
    Early Phase: In the early stages of an attack, strong defences and limited visibility are expected to constrain the agent’s progress, prompting a cautious, reconnaissance-driven approach. This phase is typically characterised by uncertainty, minimal network compromise, and broad probing, as the agent gathers information and identifies potential vulnerabilities. To capture this behaviour, our framework isolates transitions from early phases of the attack, where only a small fraction of the network has been compromised, enabling analysis of how strategies form under high uncertainty ($C_{t}<\text{Threshold}$ and $\text{Threshold}\in(0,1)$).
@@
V-B Tactical-Level (Policy) Explainability
@@
V-B1 Temporal Q-Value Evolution and Action Preference
We track the evolution of Q-values over time to analyse how the agent’s assessment of state-action pairs changes with experience. In RL settings, Q-values begin largely undifferentiated due to sparse early feedback, and progressively evolve to reflect long-term reward expectations. Our framework leverages this temporal progression to surface how tactical preferences may emerge over time, offering insight into the agent’s evolving prioritisation of specific actions. The Q-values are updated via Q-learning rule [28]:
@@ V-B1 Temporal Q-Value Evolution and Action Preference
V-B2 Policy Interpretation via Prioritised Experience Replay
@@
By integrating strategic-level progression with tactical decision analysis, our framework provides a comprehensive and temporally grounded explanation of RL-based attacker behaviour. It elucidates not only the actions taken by the agent, but also the underlying rationale, timing, and policy evolution driving those decisions. This layered interpretability enhances transparency, supports informed defensive planning, and enables the development of phase-aware countermeasures aligned with the agent’s behavioural dynamics and learning patterns. Moreover, its generalisable design allows seamless adaptation to diverse environments and attacker profiles, making it a practical tool for analysing and refining autonomous cyber operations in both research and operational contexts.
Figure 2: Cumulative rewards comparison of attacker policies across environments. The shaded region represents the standard deviation of cumulative rewards across training steps for each agent. (Results indicate that while Exploiting-DQL attains the highest rewards by leveraging a fixed, trained DQL policy, standard DQL is the most effective core agent, providing high cumulative rewards with stable and scalable learning across all environments).
VI Experimental Evaluation
We first evaluate our framework across five experimental setups designed to highlight key explainability dimensions over raw performance: (1) baseline policy selection, (2) exploration–exploitation behaviour, (3) phase-aware behavioural shifts, (4) temporal Q-value evolution, and (5) strategic learning signals via PER. These experiments are designed to reveal how and why attacker agents adapt over time, surfacing interpretable patterns such as evolving intent, phase transitions, action prioritisation, and critical learning events.
We then extend this evaluation with four additional studies (Setups 6–9) that deepen the analysis along algorithmic and behavioural dimensions, including cross-algorithm generalisability, policy-level dynamics of confidence and uncertainty, joint policy–performance explainability, and behavioural lock-in through dominant action evolution. Together, these experiments explain not only what decisions the agent makes, but how, when, and why its learning dynamics emerge, stabilise, and ultimately collapse, yielding a coherent mechanistic account of attacker behaviour beyond surface-level performance metrics.
Environment Description. We evaluated our framework using CyberBattleSim[17] across two core environments: (1) CTF, a 12-node network designed for targeted attack scenarios, and (2) CyberBattleChain, with three scalable variants, CC22, CC100, and CC500, representing networks of 22, 100, and 500 nodes, respectively. Table I summarises these configurations, where “Size” indicates the number of exploitable nodes and “Max Nodes” represents the full network, including non-exploitable components such as firewalls and auxiliary infrastructure.
TABLE I: Summary of CyberBattleSim environments, showing exploitable nodes (Size) and total nodes (Max Nodes).
@@ VI Experimental Evaluation
Figure 3: Impact of Exploration Strategies on cumulative rewards across environments (Results show that Standard Exploration achieves broader state-space coverage, yielding higher rewards and facilitating more robust policy convergence).
@@
Figure 4: Impact of Exploration Strategies on node discovery rate across environments (Results show that Standard Exploration consistently enables faster and broader discovery, facilitating more informed policy learning).
VI-A Setup 1: Baseline Policy Selection
@@
Figure 5: Cumulative rewards comparison in early vs. late attack phases across environments (Results show that agents achieve higher rewards in the late phase, indicating a progression from exploratory behaviour to exploitation-focused strategies as training advances).
Results. We trained each policy with a consistent optimisation setup, adjusting the number of episodes and iterations per episode to reflect the complexity of each environment. Specifically, the agent was trained for 20 episodes (1500 iterations) in CTF, 20 episodes (2000 iterations) in CC22, 100 episodes (2000 iterations) in CC100, and 200 episodes (4000 iterations) in CC500 environment.
Figure 2 summarises performance across environments and the results show that DQL significantly outperformed the baseline agents in terms of stability, scalability, and cumulative reward. Classical methods such as Tabular Q-Learning and Credential Lookup struggled in larger environments (CC100, with no viable results in CC500). Random search underperformed in all structured settings. While the Exploiting DQL variant achieved slightly higher cumulative rewards, it is not a standalone method; it relies on a fully trained DQL agent and simply exploits the final policy without any further training or adaptation. As such, it does not reflect the ongoing decision-making process that our explainability framework aims to explain. We therefore selected DQL as the attacker policy for all subsequent experiments. We emphasise that the results in Figure 2 serve to establish a realistic setting to demonstrate the utility of our multi-layer explainability framework.
Our benchmarking shows DQL as the most effective and scalable attacker policy across all environments. While the Exploiting DQL variant achieves marginally higher rewards, it leverages a trained DQL policy without further learning, making it less representative of adaptive behaviour. Classical approaches like Random Search, Tabular Q-Learning, and Credential Lookup fail to scale, particularly in larger networks (CC100 and CC500). We therefore adopt DQL for all subsequent experiments, ensuring that our explainability framework is evaluated on a stable and representative attacker policy.
VI-B Setup 2: Exploration-Exploitation Dynamics
To analyse how exploration-exploitation strategies shape agent behaviour, we compare two strategies: Early Exploitation and Standard Exploration and track their impact on cumulative reward and node discovery rate across environments. This reveals whether the agent behaves opportunistically, methodically, or inefficiently over time.
Results. For Early Exploitation, we initialised the exploration rate at $\epsilon=0.001$ and increased it additively by 0.005 per step until reaching $\epsilon=0.9$. In contrast, Standard Exploration began with $\epsilon=0.9$ and decayed multiplicatively to $\epsilon=0.01$ using a decay factor of 0.95. Training was conducted for 30 episodes (200 iterations per episode) in CTF and CC22, 40 episodes (300 iterations per episode) in CC100, and 100 episodes (300 iterations per episode) in CC500. Figure 3 shows the impact of exploration strategies on cumulative rewards across environments. In smaller topologies (CTF and CC22), Standard Exploration quickly outperformed Early Exploitation, benefitting from more aggressive probing. In larger environments (CC100 and CC500), this advantage widened considerably: Early Exploitation frequently stagnated at lower cumulative rewards, while Standard Exploration achieved sustained improvement and a late-stage reward surge in CC500 after broad, systematic probing. These patterns are echoed in the node discovery rates in Figure 4 where Standard Exploration uncovered more nodes earlier, while Early Exploitation remained constrained.
Our findings demonstrate that the choice of exploration strategy fundamentally drives both the efficiency of learning and the tactical profile of the agent. Standard Exploration promotes broader state-space coverage and faster reward accumulation, enabling the agent to form well-informed policies through deliberate reconnaissance and systematic probing. In contrast, Early Exploitation narrows the exploration horizon, leading to opportunistic but less comprehensive behaviour that limits performance in larger and more complex environments such as CC100 and CC500. By exposing these strategy-driven behavioural signatures, our framework provides interpretable evidence of how exploration dynamics influence policy formation, enabling rigorous analysis of attacker decision-making and implications for scalable, realistic adversarial simulations.
@@ VI-B Setup 2: Exploration-Exploitation Dynamics
Figure 6: Emergence of action preferences via state-aggregated Q-values across episodes. The highlighted dominant action marks the shift from exploration to high-value, environment-specific tactics. DQL’s sharp Q-value gradients reflect discrete exploitation bursts as the agent locks onto effective attack sequences. Action indices for CyberBattleChain (1–15) and CTF (1–18) correspond to the x-axis and are listed in Table II and Table IV, respectively.
VI-C Setup 3: Phase-Aware Behaviour Evolution
This setup supports phase-aware explainability by segmenting attacker behaviour into early and late stages using the compromise ratio $C_{t}$, defined as the proportion of compromised nodes, to highlight when strategic shifts occur.
Results. Agents were trained for 20 episodes (200 iterations) in the CTF and CC22 environments, 100 episodes (400 iterations) in CC100, and 200 episodes (400 iterations) in CC500. The early phase is defined as $C_{t}<0.5$, while the late phase corresponds to $C_{t}\geq 0.5$. Figure 5 shows that early-phase rewards remain modest due to limited visibility and cautious probing. In contrast, late-phase curves exhibit higher rewards and steeper growth, reflecting increased access, confidence, and exploitation capability.
In CyberBattleChain environments, the attack path is fixed, with each compromised node granting access to the next. Behavioural variation therefore reflects how efficiently the agent escalates privileges, selects exploits, and avoids traps, rather than differences in path selection. The observed late-phase deceleration in these environments likely reflects structural constraints rather than a fundamental strategic shift. By contrast, the CTF environment presents a more open topology, where phase-aware analysis reveals how strategies evolve across multiple possible paths, including which services and nodes are prioritised once situational awareness improves. Such insights can assist defenders in identifying high-risk assets and anticipating the timing of escalation.
The $C_{t}=0.5$ threshold serves as an interpretable behavioural inflection point rather than a hard operational rule. Below this threshold, agents predominantly engage in reconnaissance-driven exploration under high uncertainty, whereas beyond it, expanded visibility and access enable more targeted, high-impact exploitation. This phase-based segmentation provides a principled and consistent basis for analysing how attacker behaviour evolves over the course of training.
Our framework simulates an adaptive attacker that improves across training episodes via policy (parameter) updates, without using explicit episodic memory. Segmenting attacker behaviour into early and late phases using the network compromise ratio ($C_{t}$) reveals how strategies evolve during training. In open environments like CTF, agents shift from cautious probing to targeted exploitation as access expands, whereas in constrained settings such as CyberBattleChain, behaviour plateaus earlier due to structural limits on strategic diversity. This phase-aware analysis exposes when and how escalation occurs and helps defenders assess how network structure shapes adversarial learning. The resulting phase-specific reward curves also provide diagnostic value, enabling deeper evaluation of policy generalisation, convergence dynamics, and environmental impacts on learned tactics. Equivalent PPO experiments confirm these trends and the persistence of structural transitions across learning paradigms (see §VI-F).
TABLE II: Action indices and descriptions for Chain environments (CC22, CC100, CC500). These indices (1-15) correspond to the x-axis in Figure 6 and are used to interpret the emergence of action preferences.
Index Abstract Action Description
1 local(CrackKeePassX) Local attempt to crack KeepPassX and retrieve RDP credentials on Linux nodes.
2 remote(ProbeLinux) Remote probe to confirm whether a target host is running Linux.
3 local(ScanBashHistory) Local scan of bash history to discover references to other machines.
4 remote(ProbeWindows) Remote probe to determine if a target host is running Windows.
5 local(CrackKeePass) Local attempt to crack KeepPass on Windows nodes to obtain SSH credentials.
6 connect(SSH) Network connection to a target over the SSH service using known credentials.
7 connect(RDP) Network connection to a target over the RDP service using known credentials.
8 connect(SSH-key) SSH connection attempt using a leaked or cached SSH key.
9 connect(MySQL) Connection to a MySQL service on the target host.
10 connect(HTTPS) HTTPS service connection, typically for basic enumeration.
11 connect(GIT) Connection to a Git service port on the target.
12 connect(PING) ICMP ping action to check reachability of the target host.
13 local(ScanExplorerRecentFiles) Local scan of Windows Explorer recent files (trap on Linux systems).
14 local(SudoAttempt) Local attempt to escalate privileges via sudo on Linux.
15 connect(su) Connection or command using su credentials on the target host.
VI-D Setup 4: Temporal Q-Value Analysis